OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. Medal for the steps and for the correct solution x - y = 3 and x + y = 7

OpenStudy (anonymous):

if you take the two equations and add them. x - y = 3 x + y = 7 add the x's, y's and integers so x + x = 2x; (-y) + y = 0 and 3 + 7 = 10 the result is 2x = 10. x = 5 replace x into an equation for 5 - y = 3; so -y = -2, y = 2